### v1

# Superexchanges and Charge Transfer in the La3Ni2O7 Thin Films

Abstract
[…]
Introduction.- The discovery of pressure induced superconductivity in the Ruddlesden-Popper (RP) nickelates - first in bilayer La3Ni2O7 ($T_{c}\approx$ 80 K) [1] and subsequently trilayer La4Ni3O10 ($T_{c}=20\sim 30$ K) [2, 3], - has spurred extensive research interest. This has led to both numerous theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and extensive experimental studies [3, 35, 36, 37, 38, 39, 40, 41] .A crucial next step for the in-depth understanding and practical application of nickelate superconductors is achieving superconductivity at ambient pressure. Current proposed strategies to reach this milestone focus mainly on chemical substitution [42, 43, 44]and the application of strain [45, 46, 47, 48].
Recent experiments report superconductivity at ambient pressure with $T_{c}$ exceeding 40 K in La3Ni2O7 (LNO) [49] and (La,Pr)3Ni2O7 [50] thin films. These advancements circumvent the technical challenges of high-pressure experiments, while enabling advanced characterization techniques, such as angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling microscopy (STM). Superconductivity in thin films is achieved by growing LNO films on substrates with reduced lattice constants (such as LaAlO3 (LAO) or SrLaAlO4 (SLAO)), which imposes significant in-plane biaxial compressive strain on the LNO film samples [51]. The compressive strain forces the crystal structure of the thin film to exhibit tetragonal symmetry, mimicking the Fmmm or I4/mmm phase observed in bulk materials under high pressure, thus effectively suppressing the octahedral tilting observed in the Amam phase of bulk materials under ambient pressure [52, 45, 53]. This observation suggests that strain could decouple and independently control in-plane and out-of-plane electronic and magnetic properties, providing a unique avenue for optimizing superconductivity at ambient pressure. Although strain-induced thin films have successfully reproduced the highly symmetric structure and superconductivity of high-pressure bulk materials, a fundamental assumption on the similiary of their low-energy effective Hamiltonians remains untested. This raises a key question about how structural changes in the LNO film affects the strength of magnetic superexchange couplings and the distribution of the charge carriers. These two factors are believed to be directly linked to the system’s superconducting properties.
To address above question, here we study an 11-band Hubbard model that includes four $3d_{x^{2}-y^{2}}$ / $3d_{3z^{2}-r^{2}}$ orbitals of nickel, and seven most relevant $2p$ orbitals of oxygen in the NiO2 bilayer per site [54]. We perform numerical exact determinant quantum Monte Carlo (DQMC) simulations [55, 56] and cellular dynamical mean-field theory (CDMFT) calculations [57, 58] in the normal state of the system. Our result suggests that the LNO thin film exhibits weaker inter-layer (IT) couplings, while the strength of the intra-layer (IR) antiferromagnetic (AFM) couplings remain nearly unchanged comparing to the bulk LNO. We study the charge-transfer properties [59] in the Zaanen-Sawatzky-Allen (ZSA) scheme [60], which is crucial for understanding $T_{c}$ in cuprate superconductors [61, 62, 63, 64, 65]. We reveal that holes introduced by doping the pristine system with heterovalent elements are distributed approximately equally between the in-plane and out-of -plane orbitals, whereas electron doping leads to charge carriers distributed with a 3:1 ratio. By comparing with the doping phase diagram [66] of superconducting La3-xSrxNi2O7 films, we argue that the occupation number of $d_{3z^{2}-r^{2}}$ is crucial for determining the type of the superconductivity [67], namely, whether from the Hund’s coupling $J_{H}$ [21, 20, 19] or the hybridization $V$ between $e_{g}$ orbitals [68].
Figure 1: Schematic of the 11-band bilayer LNO thin film lattice model with the seven major hopping parameters and the illustration of the two-dimensional NiO2 lattice for DMQC or CDMFT simulations. (a). The orange, blue and green electron clouds denote $d_{3z^{2}-r^{2}}, Ni-3$$d_{x^{2}-y^{2}}$ and O-2pp orbitals. For brevity, some of the orbitals on plane $b$ is not displayed. Four hopping processes between Ni-$d$ and O-$p$ orbitals predominantly contribute to superexchange interactions between Ni-$d$ orbitals are: $t_{1}=-1.4795,t_{2}=0.6605,t_{3}=-1.4565,t_{6}=1.2045$. The site energies are: $\varepsilon_{d_{x^{2}-y^{2}}}=-0.962,\varepsilon_{d_{3z^{2}-r^{2}}}=-1.070,\varepsilon_{p_{x}/p_{y}}=-4.650,\varepsilon_{p_{z}}=-4.078,\varepsilon_{p_{z}^{{}^{\prime}}}=\varepsilon_{p_{z}^{{}^{\prime\prime}}}=-2.977$ [54]. The superexchanges between the intra-layer $d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals (not shown here) vanishes due to symmetry. Here $t_{4}=0.5675,t_{5}=0.4315,t_{7}=0.4115$ denote electrons hopping processes between O-$p$ orbitals. (b). Dark balls denote nickel positions while light balls show oxygen positions in the NiO2 plane. The 6 $\times$ 6 unit cells used in DQMC simulations is shown here directly, while the 2 $\times$ 2 CDMFT effective cluster is outlined by a blue dashed rectangle. For clarity, only one NiO2 layer is shown here.
Model and method.- To study the superexchange couplings in LNO films, we consider a high-energy effective model including eleven Ni and O orbitals. The Halmiltonian is written as:
$H=$ $\sum_{ijab\alpha\beta\sigma}t_{ia\alpha,jb\beta}c_{ia\alpha\sigma}^{\dagger}c_{jb\beta\sigma}+\sum_{a\alpha\sigma}(\varepsilon_{a\alpha}-\mu)n_{a\alpha\sigma}-\sum_{ia\alpha\sigma}E_{\mathrm{DC}}^{a\alpha}n_{ia\alpha\sigma}^{d}$ (1)
$+U\sum_{ia\alpha}n_{ia\alpha\uparrow}^{d}n_{ia\alpha\downarrow}^{d}+\sum_{iab\alpha\beta\sigma\sigma^{\prime}}(V-J_{H}\delta_{\sigma\sigma^{\prime}})n_{ia\alpha\sigma}^{d}n_{jb\beta\sigma^{\prime}}^{d},$
where i/j, a/b, $\alpha/\beta$, labels lattice sites, NiO2 layers and orbitals respectively, $\sigma$ denotes spin degree of freedom. The superscript $d$ indicates that only Hubbard-$U$ in Ni-$3d$ orbitals is considered. Here $t_{ijab\alpha\beta}$ denotes electrons hoppings. $\varepsilon_{a\alpha}$ stands for site-energy of $\alpha$ orbital on $a$ layer. $U$ is the Hubbard repulsion between two electrons on the same $d$ orbital, and $V=U-2J_{H}$ is for that on two different orbitals [69], where $J_{H}$ is the Ising-type Hund’s coupling [15]. Here, pair-hopping and spin-flip terms in Hund’s coupling are neglected. The double-counting term $E_{\mathrm{DC}}$ is subtracted in both the DQMC and CDMFT simulations. We adopt Held’s formula [70] for the double-counting, $E_{DC}=\frac{1}{3}(U+2V-J_{H})(n^{0}_{d}-0.5)$ where $n_{d}^{0}$ is the occupation number of $d$-orbital in the non-interacting limit. In our case $n_{d}^{0}\approx 2.14$. We adopted hopping parameters and site-energies from a DFT calculation on the double-stacked LNO thin film in Ref. [54]. See also Figure 1 (a) and its caption for all the hopping parameters and site-energy values. Based on DFT results [54], we set the chemical potential $\mu=0$ in the Hamiltonian to represent the pristine single-crystalline LNO thin films under ambient pressure. For the purpose of exploring different doping regimes, we adjust $\mu$ and define the hole concentration $n_{h}=\frac{1}{4}(22-\sum_{a\alpha\sigma}n_{a\alpha\sigma})$, representing the average number of holes per $d$-orbital per site. Unless stated otherwise, we use $U=7$ eV as default parameters [15, 59] for Ni $e_{g}$ orbitals. For the calculations on bulk LNO, we employ a set of parameters for the 11-bands model [4, 15].
We work on a 2D square lattice containing up to 6 $\times$ 6 $\times$ 11 = 396 orbitals for DQMC simulation and an effective impurity model with 2 $\times$ 2 $\times$ 11 = 44 orbitals for CDMFT study, as shown in Figure 1 (b). Due to the fermionic sign problem in DQMC, our simulations are restricted to the high-temperature regime, where the lowest temperature $T=0.25$ (in eV units) is achieved with an average sign $\langle s\rangle\approx 0.79$. The CDMFT is employed as a complementary method accessing low-temperature regimes, where $T\sim 0.08$ can be reached while maintaining an average sign $\langle s\rangle\approx 0.56$ in the impurity solver.
Superexchanges.- We now discuss the magnetic exchange couplings in LNO thin films, comparing them with those in bulk LNO. As demonstrated in Figure 1 (a), four hopping processes ($t_{1}$, $t_{2}$, $t_{3}$ and $t_{6}$) contribute to the main superexchange interactions, with the corresponding values of $J_{\parallel}\sim 0.084$ eV and $J_{\perp}\sim 0.135$ eV, as summarized in Table 1. In the atomic limit, the charge-transfer picture for spin-1/2 electron system leads to an energy difference between the spin singlet and spin triplet composed of the two Ni $d$-orbitals,
$J$ $=E_{\uparrow\downarrow}-E_{\uparrow\uparrow}$
$=-2|t_{pd_{1}}t_{pd_{2}}|^{2}\left(\frac{1}{U+\Delta_{pd_{2}}-\Delta_{pd_{1}}}\left(\frac{1}{U+\Delta_{pd_{1}}}\right)^{2}\right$.
$\quad\left.{}+\frac{1}{U+\Delta_{pd_{1}}-\Delta_{pd_{2}}}\left(\frac{1}{U+\Delta_{pd_{2}}}\right)^{2}\right$.
$\quad\left.{}+\frac{1}{2U+\Delta_{pd_{1}}+\Delta_{pd_{2}}}\left(\frac{1}{U+\Delta_{pd_{1}}}+\frac{1}{U+\Delta_{pd_{2}}}\right)^{2}\right),$ (2)
the $t_{pd_{1,2}}$ represent the hopping between the Ni-$d_{1,2}$ orbitals and the O-$p$ orbital, and the energy gap between the lower Hubbard band of the $d$-orbitals and the O-$p$ orbital is given by: $\Delta_{pd_{1,2}}=\varepsilon_{d_{1,2}}-\varepsilon_{p}$, Note that here the double-counting term is already include in $\varepsilon_{d_{1,2}}$. Equation 2 presents the second-order approximation result from perturbation theory, which describes the superexchange between two neighboring Ni-$3d$ orbitals in the atomic limit. Our result of $J$ for various $U$ can be found in Table 1. One can see that as $U$ increases, the AFM coupling coefficient $J$ decreases. For example, at $U=3$ eV, $J_{\perp}=0.325$ eV, and at $U=7$ eV, $J_{\perp}=0.135$ eV.
Figure 2: The spin-spin correlation function $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ for four neighboring $d$-orbitals is shown in numbers to demonstrate the relative strength of the antiferromagnetic superexchange couplings in the system. The orange and blue symbols represent $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals respectively. We have used the specific values $U=7$ and $J=0.15U$, very similar to those obtained from constrained RPA [14]. (a) are from DQMC at half-filling ($n_{h}=1$, $\mu=1.5$) at $T=0.25$, and (b) are from CDMFT at $\mu=0$ ($n_{h}\approx 1.234$). Note that the magnetic correlations between the on-site inter-layer $d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals, which are due to Hund’s coupling, are not shown here.
Table 1: The value of superexchanges $J$ at various Hubbard $U$ from Equation 2. The vertical inter-layer AFM coupling $J_{\perp}$, intra-layer AFM coupling $J_{\parallel}$ and their ratio $J_{\parallel}/J_{\perp}$ are shown. The double counting term used here is the same as in the DQMC and CDMFT simulations: $E_{DC}=0.473U$ and $E_{DC}=0.589U$ for $d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals, respectively. The calculations are performed up to sixth-order precision.
  UU/eV   $J_{\perp}$/eV   $J_{\parallel}$/eV   $J_{\parallel}/J_{\perp}$
  3   0.325   0.243   74.6%
  4   0.263   0.181   68.6%
  5   0.207   0.137   66.1%
  6   0.165   0.106   64.5%
  7   0.135   0.084   62.3%
In DQMC calculations, the magnetic correlation functions can be exactly measured at high temperature. In Figure 2 (a) and (b), we show the magnitudes of the spin correlation function $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ between neighboring $d$- orbitals at $T=0.25$ and half-filling from DQMC, and $\mu=0$ at $T=0.125$ from CDMFT, respectively. Here $S_{ia\alpha}$ is spin operator with site index $i$, layer $a$ and orbital $\alpha$. This results depicts the relative strengths of the major exchange couplings in the LNO thin film. In Figure 2 (a), at $T=0.25$ and half-filling, comparing to the bulk LNO, the dominant AFM exchange coupling is also the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling, with a value of $\langle S\cdot S\rangle=-0.057(2)$. The next strongest AFM exchange comes from IR nearest-neighboring (NN) $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ orbitals, with $\langle S\cdot S\rangle=-0.046(7)$. Relative to bulk, this value increases significantly in thin films, reaching 81% of AFM exchange between IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ orbitals. The same ratio in bulk LNO it is only 60% [15]. Moreover, the strengths of IR NN $d_{x^{2}-y^{2}}-d_{3z^{2}-r^{2}}$ and $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ AFM correlations are $\langle S\cdot S\rangle=-0.024(1)$ and $\langle S\cdot S\rangle=-0.017(1)$, respectively. These couplings play a minor role due to a weaker hopping between $d_{3z^{2}-r^{2}}$ and $p_{x}/p_{y}$ orbitals in films. Notably, the IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ exchange coupling within a unit cell is the weakest of these. In Figure 2 (b), the dominant AFM exchange coupling is also the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling from CDMFT approaches when $T=0.125$ and $\mu=0$. Notably, the strength of the correlation between IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ orbitals is the weakest of the five magnetic couplings. In essence, it is clear that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlation dominates at low temperatures, suggesting that $d_{3z^{2}-r^{2}}$ orbitals are the key components responsible for strong-correlation-driven Superconductivity.
Figure 3: $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ between pairs of nearest-neighboring (NN) d-orbitals as a function of temperature $T$ at pristine configuration, as obtained from DQMC simulations. The squares indicate results for the intra-layer (IR) $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ correlations, while the circles show the results for the inter-layer (IT) $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlations.
Figure 3 shows the temperature dependence of AFM correlations $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ obtained from DQMC simulations at $\mu=0$. Compared with the 11-band model of bulk LNO. It is observed that the AFM correlations of bulk LNO are stronger than those in the thin film, for both IR and IT interactions. This could provide an explanation for the reduced transition temperature $T_{c}$ in the thin film, caused by the weaker AFM superexchange compared to bulk LNO. In film, the reduction in the magnetic correlations between the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ orbitals is much more prominent than the reduction in IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component. The latter is almost indistinguishable between film and bulk. Furthermore, it can be seen that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ magnetic correlation in thin film is not overly stronger than that of IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component, contrasting with the situation in bulk, where the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlation is much stronger than the IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ correlations [15]. This results imply that both the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ component and the IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component of AFM correlations may play crucial roles in the observed superconductivity in the LNO thin film at ambient pressure.
Figure 4: Charge transfer properties and charge carrier evolution of the La3Ni2O7 thin film. (a). DQMC and CDMFT result of hole concentration $n_{h}$ as a function of hole chemical potential $\mu_{h}$. An inflection point suggests that a charge transfer gap opens at half-filling when the hole chemical potential $\mu_{h}$ approaches $\mu_{h}\sim-1.55$. (b). Carrier concentration as a function of Hund’s coupling $J_{H}$ for each orbital in the LNO thin film under half-filling (dashed line) and pristine configuration ($\mu=0$, solid line) conditions by CDMFT at $T=0.125$. The left axis represents the in-plane and out-of-plane hole concentrations, while the right axis shows the hole concentration of oxygen orbitals. (c) and (d) show electron(hole) concentration variation $\Delta n_{e(h)}^{\alpha}$ as functions of chemical potential $\mu_{e(h)}$ and $\text{Sr}^{2+}(\text{Ce}^{4+})$ nominal doping levels $x$, respectively. Here, $\Delta n_{h}^{\mathrm{IP}}=\Delta n_{h}^{d_{x^{2}-y^{2}}}+\Delta n_{h}^{p_{x}}+\Delta n_{h}^{p_{y}}$ counts the number of holes in the in-plane $d_{x^{2}-y^{2}}$ orbital and $p_{x}/p_{y}$ orbitals combined, while $\Delta n_{h}^{\mathrm{OP}}=\Delta n_{h}^{d_{3z^{2}-r^{2}}}+0.5(\Delta n_{h}^{p_{z}}+\Delta n_{h}^{p_{z}^{\prime}})$ counts the number of holes in the out-of-plane. Note that the $\Delta n_{h}^{\alpha}$ here represents the difference in hole concentration relative to the pristine LNO thin film (Ni-$3d^{7.5}$), i.e. Hole concentration variation $\Delta n_{h}^{\alpha}\equiv n_{h}(\mu)-n_{h}(\mu_{h}=0)$. When Sr doping level $x$ = 0.203, the nominal hole content in La3-xSrxNi2O7 corresponds to $\mu_{h}=0.2$, while at $\mu_{h}$ of 0.4, the equivalent Sr doping level is $x$ = 0.406. At this doping level, superconductivity is no longer observed according to Ref. [66]. (e). The electron concentration variation $\Delta n_{e}^{\alpha}$ as a function of the electron chemical potential $\mu_{e}$ in the regime where $\mu_{e}>1.55$ (corresponding to electron doping with $n_{e}>1$) and hole concentration variation $\Delta n_{h}^{\alpha}$ in the regime where hole doping with $n_{h}=2-n_{e}>1$($n_{e/h}=\sum_{a,\alpha}n^{a\alpha}_{e/h}$), respectively. Here, $\alpha$ denotes distinct $\alpha$-orbitals and charge concentration variation $\Delta n_{e/h}^{\alpha}\equiv n_{e/h}(\mu)-n_{e/h}(\mu_{e}=1.55)$. The results for (c), (d) and (e) above are from DQMC simulations at $T=0.3$.
Charge transfer properties.- We now focus on the charge transfer properties in 11-band Hubbard model, Figure 4 illustrates the vary of the hole/electron concentration changes in different orbitals as functions of chemical potentials $\mu$ and Hund’s coupling $J_{H}$. We observe an inflection point at $n_{h}=1$ ($\mu_{h}\sim-1.55$) with an almost flat compressibility ($\partial n_{h}/\partial\mu_{h}$) in Figure 4 (a), which implies there could an opening charge gap at half-filling. Note that here due to the escalated temperature we use, the $\mu-n_{h}$ curve is not fully flat, denoting a not vanishing compressibility. Notably, we hypothesize that this charge gap is smaller compared to that of bulk LNO due to the significant impact of chemical potential on the charge carrier concentration.
We also present doping-induced, orbital-resolved distribution of doped charge carriers in Figure 4 (e). As the hole chemical potential $\mu_{h}\equiv-\mu$ and electron chemical potential $\mu_{e}\equiv\mu$ are tuned from half-filling ($\mu_{h}=-\mu_{e}\approx-1.55$), holes/electrons to populate distinct Ni-$d$ and O-$p$ orbitals. This process, quantified by $\Delta n_{h/e}^{\alpha}$, is shown in Figure 4 (e). Consistent with bulk LNO [15, 14, 71], doped holes predominantly occupy O-$p$ orbitals (Right of Figure 4 (e)), while doped electrons primarily occupy Ni-$d$ orbitals (Left of Figure 4 (e)). This confirms the charge-transfer nature of the insulating state in this system at the half-filling state[60]. Moreover, we observe that the $d_{x^{2}-y^{2}}$ orbital contain a higher proportion of doped holes than the $d_{3z^{2}-r^{2}}$ orbital in the case of hole doping ($\mu_{h}>-1.55$), as seen in Figure 4 (e). This reflects the stronger correlation of the $d_{3z^{2}-r^{2}}$ orbital compared to the $d_{x^{2}-y^{2}}$ orbital under hole doping, as evidenced by the significant impairment of superexchange coupling in cuprates caused by hole doping [62, 61].
To shed light on the effective low-energy model study that neglect oxygen degree of freedom, we group the Ni-$3d$ and O-$p$ orbitals into two classes: the out-of-plane (OP) correlated orbitals, including $d_{3z^{2}-r^{2}}$, $p_{z}$ and $p_{z}^{\prime}$ and the in-plane (IP) correlated orbitals, comprising $d_{x^{2}-y^{2}}$ and $p_{x}/p_{y}$ orbitals. And analyze their carrier concentration under hole/electron doping. The results, shown in Figure 4 (d), indicated that the additional holes doped into the LNO film are distributed nearly equally between the IP and OP orbitals, with a slight excess of holes in the IP orbitals. In the case of electron doping, the ratio of doped electrons in the IP to the OP orbitals is approximately 3:1 until the system reaches half-filling. This contrasts with the 1:1 ratio observed in hole doping, which suggests an asymmetric doping-dependent superconducting phase diagram. In experiments, Sr-doping introduce extra holes into the system [66], while possible Ce-doping can introduce extra electrons into the pristine system.
In comparison to bulk LNO, we find that under the same chemical potential variation, the doping variable of LNO film is approximately twice that of bulk LNO [15]. This suggests that LNO films exhibit enhanced charge transfer capabilities compared to the bulk LNO under in-plane biaxial compression, meaning the LNO thin film possesses a smaller charge transfer gap $\Delta$ [72]. This finding aligns with recent spectroscopic experimental results [73], which propose that two types of Zhang–Rice singlets exist in the LNO film and that the fundamental classification of in-plane charge transfer properties is identical to that of the cuprate superconductors.
Furthermore, it is discovered that the presence of Hund’s coupling $J_{H}$ in LNO films under half-filling strongly suppresses orbital polarization, thereby stabilizing a high-spin $S=1$ state. This phenomenon is illustrated by the gray and red dashed lines in Figure 4 (b), which correspond to the Hund’s first law. In contrast, in the pristine state, the $J_{H}$ coupling is significantly localized, and holes still preferentially occupy the in-plane orbitals, as indicated by the gray and red solid lines in Figure 4 (b).
Discussion and conclusion.- The observation of superconductivity in La3Ni2O7 thin films at ambient pressure [49, 50] has spurred proposals of several effective interacting models to investigate the pairing symmetry [74, 75, 51, 52], building on insights from ab initio calculations [47, 48, 54, 74, 76]. These effective models primarily focus on parameterizing the magnetic correlations between Ni $d$-orbitals through direct exchange couplings, while the contribution of oxygen degrees of freedom is neglected. In contrast our 11-band Hubbard model, which incorporates both Ni-$3d$ and O-$2p$ orbitals, provides a more comprehensive picture by quantifying the relative strengths of superexchange couplings between distinct Ni-$3d$ orbitals. Crucially, we identify two spin-singlet bands exhibiting markedly different hole concentrations ($n_{h}$) and correlation strengths. We observe the significant reduction of IT magnetic exchange coupling, which makes the IR and IT coupling strengths comparable, in sharp contrast to the IT-dominated bulk material. Furthermore, we observe a pronounced asymmetry in charge distribution upon doping. It should be noted that some studies suggest Hund’s coupling $J_{H}$ can transfer antiferromagnetic correlation from $d_{3z^{2}-r^{2}}$ orbital to $d_{x^{2}-y^{2}}$ orbital [21, 20, 19, 77]. Based on our results, we find that a large $J_{H}$ does indeed enhance the inter-layer $d_{x^{2}-y^{2}}$ correlation. However, the large $J_{H}$ does not significantly impact the carriers distribution in the $e_{g}$ orbitals. Our analysis of both spin-spin correlations and charge-transfer properties consistently indicates that the inter-layer $Ni-3d_{3z^{2}-r^{2}}$ orbital antiferromagnetic exchange interaction remains an important factor that primarily determines the nature of the bilayer nickelate superconducting pairing. And the contribution of more itinerant Ni-$d_{x^{2}-y^{2}}$ orbital in to both strange-metal (SM) and superconducting (SC) states remains complex and requires further investigations.
Acknowledgements.- We thank the useful discussion with Zhihui Luo and Wenyuan Qiu. This project was supported by NSFC-12494591, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).

### v2

@@
 […]
-Introduction.- The discovery of pressure induced superconductivity in the Ruddlesden-Popper (RP) nickelates - first in bilayer La3Ni2O7 ($T_{c}\approx$ 80 K) [1] and subsequently trilayer La4Ni3O10 ($T_{c}=20\sim 30$ K) [2, 3], - has spurred extensive research interest.
-This has led to both numerous theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and extensive experimental studies [3, 35, 36, 37, 38, 39, 40, 41] .A crucial next step for the in-depth understanding and practical application of nickelate superconductors is achieving superconductivity at ambient pressure.
-Current proposed strategies to reach this milestone focus mainly on chemical substitution [42, 43, 44]and the application of strain [45, 46, 47, 48].
-Recent experiments report superconductivity at ambient pressure with $T_{c}$ exceeding 40 K in La3Ni2O7 (LNO) [49] and (La,Pr)3Ni2O7 [50] thin films.
-These advancements circumvent the technical challenges of high-pressure experiments, while enabling advanced characterization techniques, such as angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling microscopy (STM).
-Superconductivity in thin films is achieved by growing LNO films on substrates with reduced lattice constants (such as LaAlO3 (LAO) or SrLaAlO4 (SLAO)), which imposes significant in-plane biaxial compressive strain on the LNO film samples [51].
-The compressive strain forces the crystal structure of the thin film to exhibit tetragonal symmetry, mimicking the Fmmm or I4/mmm phase observed in bulk materials under high pressure, thus effectively suppressing the octahedral tilting observed in the Amam phase of bulk materials under ambient pressure [52, 45, 53].
-This observation suggests that strain could decouple and independently control in-plane and out-of-plane electronic and magnetic properties, providing a unique avenue for optimizing superconductivity at ambient pressure.
-Although strain-induced thin films have successfully reproduced the highly symmetric structure and superconductivity of high-pressure bulk materials, a fundamental assumption on the similiary of their low-energy effective Hamiltonians remains untested.
-This raises a key question about how structural changes in the LNO film affects the strength of magnetic superexchange couplings and the distribution of the charge carriers.
-These two factors are believed to be directly linked to the system’s superconducting properties.
-To address above question, here we study an 11-band Hubbard model that includes four $3d_{x^{2}-y^{2}}$ / $3d_{3z^{2}-r^{2}}$ orbitals of nickel, and seven most relevant $2p$ orbitals of oxygen in the NiO2 bilayer per site [54].
-We perform numerical exact determinant quantum Monte Carlo (DQMC) simulations [55, 56] and cellular dynamical mean-field theory (CDMFT) calculations [57, 58] in the normal state of the system.
-Our result suggests that the LNO thin film exhibits weaker inter-layer (IT) couplings, while the strength of the intra-layer (IR) antiferromagnetic (AFM) couplings remain nearly unchanged comparing to the bulk LNO.
-We study the charge-transfer properties [59] in the Zaanen-Sawatzky-Allen (ZSA) scheme [60], which is crucial for understanding $T_{c}$ in cuprate superconductors [61, 62, 63, 64, 65].
-We reveal that holes introduced by doping the pristine system with heterovalent elements are distributed approximately equally between the in-plane and out-of -plane orbitals, whereas electron doping leads to charge carriers distributed with a 3:1 ratio.
-By comparing with the doping phase diagram [66] of superconducting La3-xSrxNi2O7 films, we argue that the occupation number of $d_{3z^{2}-r^{2}}$ is crucial for determining the type of the superconductivity [67], namely, whether from the Hund’s coupling $J_{H}$ [21, 20, 19] or the hybridization $V$ between $e_{g}$ orbitals [68].
+Introduction.- The discovery of pressure-induced superconductivity in Ruddlesden-Popper (RP) nickelates, first in bilayer La3Ni2O7 ($T_{c}\approx$ 80 K) [1] and subsequently in trilayer La4Ni3O10 ($T_{c}=20\sim 30$ K) [2, 3], has stimulated extensive research.
+After numerous theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] and experimental studies [3, 36, 37, 38, 39, 40, 41, 42], various protocols have been proposed to obtain superconductivity without external pressure, including both chemical substitution [43, 44, 45, 46]and strain engineering [47, 48, 49, 50].
+Recent experiments have achieved this goal in thin films of La3Ni2O7 (LNO) [51] and (La,Pr)3Ni2O7 [52], reporting superconductivity at ambient pressure with $T_{c}$ exceeding 40K.
+This advance overcomes the experimental difficulties associated with high-pressure measurements, facilitating the use of sophisticated probes like the angle-resolved photoemission spectroscopy (ARPES) or scanning tunneling microscopy (STM).
+Superconductivity in these experiments is achieved by growing LNO films on substrates with reduced lattice constants (such as LaAlO3 (LAO) or SrLaAlO4 (SLAO)), which imposes a strong in-plane biaxial compressive strain [53].
+This strain can suppress the octahedral tilting in the Amam phase of bulk materials at ambient pressure  [54, 47, 55], stabilizing a tetragonal structure in films analogous to the high-pressure Fmmm or I4/mmm phase of superconducting bulk materials.
+Given this structural similarity, the low-energy physics is often assumed to be comparable between high-pressure bulk materials and ambient-pressure films.
+Yet, whether their effective Hamiltonians are truly equivalent is untested.
+A central unresolved issue, therefore, is to what extent the modified lattice parameters in the films alter the strength of magnetic superexchange couplings and the orbital distribution of the charge carriers.
+These two factors are believed to govern the system’s superconducting properties.
+To address above question, we study an 11-band dpd-p Hubbard model that includes four $3d_{x^{2}-y^{2}}$ / $3d_{3z^{2}-r^{2}}$ orbitals of nickel and seven most relevant $2p$ orbitals of oxygen in the NiO2 bilayer per unit cell [56].
+We perform numerically exact determinant quantum Monte Carlo (DQMC) simulations [57, 58] and cellular dynamical mean-field theory (CDMFT) calculations [59, 60] in the normal state of the system.
+Our results suggest that LNO thin film system exhibits weaker inter-layer (IT) antiferromagnetic (AFM) couplings between $d_{3z^{2}-r^{2}}$ orbitals, while the strength of the intra-layer (IR) AFM couplings between $d_{x^{2}-y^{2}}$ orbitals remain nearly unchanged comparing to the bulk LNO.
+Furthermore, we analyze the charge-transfer properties [61] in the Zaanen-Sawatzky-Allen (ZSA) scheme [62], which has been shown crucial for understanding $T_{c}$ in cuprate superconductors [63, 64, 65, 66, 67].
+We reveal that the holes introduced by doping the pristine system with heterovalent elements are distributed approximately evenly between the in-plane and out-of-plane orbitals, whereas for electron doping, it favors a roughly 3:1 occupancy ratio.
+By comparing with the experimental doping phase diagram [68] of superconducting La3-xSrxNi2O7 films, we argue that the occupation number of $d_{3z^{2}-r^{2}}$ is crucial for determining the type of superconductivity [69], namely, whether it is driven by Hund’s coupling $J_{H}$ [21, 20, 19, 70] or governed by hybridization $V$ between $e_{g}$ orbitals [71].
 Figure 1: Schematic of the 11-band bilayer LNO thin film lattice model with the seven major hopping parameters and the illustration of the two-dimensional NiO2 lattice for DMQC or CDMFT simulations.
@@
 For clarity, only one NiO2 layer is shown here.
-Model and method.- To study the superexchange couplings in LNO films, we consider a high-energy effective model including eleven Ni and O orbitals. The Halmiltonian is written as:
+Model and method.- To investigate the superexchange couplings in LNO films, we consider a high-energy effective model including eleven Ni-$d$ and O-$p$ orbitals. The Halmiltonian is given as:
 $H=$ $\sum_{ijab\alpha\beta\sigma}t_{ia\alpha,jb\beta}c_{ia\alpha\sigma}^{\dagger}c_{jb\beta\sigma}+\sum_{a\alpha\sigma}(\varepsilon_{a\alpha}-\mu)n_{a\alpha\sigma}-\sum_{ia\alpha\sigma}E_{\mathrm{DC}}^{a\alpha}n_{ia\alpha\sigma}^{d}$ (1)
@@
 where i/j, a/b, $\alpha/\beta$, labels lattice sites, NiO2 layers and orbitals respectively, $\sigma$ denotes spin degree of freedom.
-The superscript $d$ indicates that only Hubbard-$U$ in Ni-$3d$ orbitals is considered.
-Here $t_{ijab\alpha\beta}$ denotes electrons hoppings.
-$\varepsilon_{a\alpha}$ stands for site-energy of $\alpha$ orbital on $a$ layer.
+The superscript $d$ indicates that the Hubbard-$U$ is applied only to Ni-$3d$ orbitals.
+Here, $t_{ijab\alpha\beta}$ are electron hopping integrals, $\varepsilon_{a\alpha}$ stands for on-site energy of orbital $\alpha$ in layer $a$.
 $U$ is the Hubbard repulsion between two electrons on the same $d$ orbital, and $V=U-2J_{H}$ is for that on two different orbitals [72], where $J_{H}$ is the Ising-type Hund’s coupling [15].
-Here, pair-hopping and spin-flip terms in Hund’s coupling are neglected.
+Here, the pair-hopping and spin-flip terms in Hund’s coupling are neglected.
 The double-counting term $E_{\mathrm{DC}}$ is subtracted in both the DQMC and CDMFT simulations.
@@
 In our case $n_{d}^{0}\approx 2.14$.
-We adopted hopping parameters and site-energies from a DFT calculation on the double-stacked LNO thin film in Ref.
+The hopping parameters and on-site energies are adopted from a DFT calculation on the double-stacked LNO thin film in Ref.
 [56].
-See also Figure 1 (a) and its caption for all the hopping parameters and site-energy values.
-Based on DFT results [54], we set the chemical potential $\mu=0$ in the Hamiltonian to represent the pristine single-crystalline LNO thin films under ambient pressure.
+See also Figure 1 (a) and its caption for all the hopping parameters and on-site energies.
+Following DFT results [56], we set the chemical potential $\mu=0$ in the Hamiltonian to represent the pristine LNO thin films at ambient pressure.
 For the purpose of exploring different doping regimes, we adjust $\mu$ and define the hole concentration $n_{h}=\frac{1}{4}(22-\sum_{a\alpha\sigma}n_{a\alpha\sigma})$, representing the average number of holes per $d$-orbital per site.
-Unless stated otherwise, we use $U=7$ eV as default parameters [15, 59] for Ni $e_{g}$ orbitals.
-For the calculations on bulk LNO, we employ a set of parameters for the 11-bands model [4, 15].
+Unless stated otherwise, we use $U=7$ eV as our default parameters [15, 61] for Ni $e_{g}$ orbitals.
+For the comparative calculations on bulk LNO, we employ a set of parameters for the 11-bands model [4, 15].
 We work on a 2D square lattice containing up to 6 $\times$ 6 $\times$ 11 = 396 orbitals for DQMC simulation and an effective impurity model with 2 $\times$ 2 $\times$ 11 = 44 orbitals for CDMFT study, as shown in Figure 1 (b).
-Due to the fermionic sign problem in DQMC, our simulations are restricted to the high-temperature regime, where the lowest temperature $T=0.25$ (in eV units) is achieved with an average sign $\langle s\rangle\approx 0.79$.
-The CDMFT is employed as a complementary method accessing low-temperature regimes, where $T\sim 0.08$ can be reached while maintaining an average sign $\langle s\rangle\approx 0.56$ in the impurity solver.
-Superexchanges.- We now discuss the magnetic exchange couplings in LNO thin films, comparing them with those in bulk LNO.
-As demonstrated in Figure 1 (a), four hopping processes ($t_{1}$, $t_{2}$, $t_{3}$ and $t_{6}$) contribute to the main superexchange interactions, with the corresponding values of $J_{\parallel}\sim 0.084$ eV and $J_{\perp}\sim 0.135$ eV, as summarized in Table 1.
-In the atomic limit, the charge-transfer picture for spin-1/2 electron system leads to an energy difference between the spin singlet and spin triplet composed of the two Ni $d$-orbitals,
-$J$ $=E_{\uparrow\downarrow}-E_{\uparrow\uparrow}$
-$=-2|t_{pd_{1}}t_{pd_{2}}|^{2}\left(\frac{1}{U+\Delta_{pd_{2}}-\Delta_{pd_{1}}}\left(\frac{1}{U+\Delta_{pd_{1}}}\right)^{2}\right$.
+Due to the fermionic sign problem in DQMC, our simulations are restricted to a high-temperature regime, where the lowest temperature $T=0.25$ (in eV units) is achieved with an average sign $\langle s\rangle\approx 0.79$.
+The CDMFT is employed as a complementary method accessing low-temperature regime, where $T\sim 0.08$ can be reached while maintaining an average sign $\langle s\rangle\approx 0.56$ in the impurity solver.
+Superexchanges.- We first discuss the magnetic exchange couplings in LNO thin films and compare them with bulk LNO.
+As demonstrated in Figure 1 (a), four hopping processes ($t_{1}$, $t_{2}$, $t_{3}$ and $t_{6}$) contribute to the main superexchange couplings, with the corresponding coupling constants estimated as $J_{\parallel}\sim 0.084$ eV and $J_{\perp}\sim 0.135$ eV at $U=7eV$ in the atomic limit, see Table 1.
+In the atomic limit, the charge-transfer picture for spin-1/2 electron system suggests that the energy difference between the spin singlet and spin triplet states composed of the two Ni $d$-orbitals can be written as,
+$J$ $=E_{\uparrow\uparrow}-E_{\uparrow\downarrow}$
+$=2|t_{pd_{1}}t_{pd_{2}}|^{2}\left(\frac{1}{U+\Delta_{pd_{2}}-\Delta_{pd_{1}}}\left(\frac{1}{U+\Delta_{pd_{1}}}\right)^{2}\right$.
 $\quad\left.{}+\frac{1}{U+\Delta_{pd_{1}}-\Delta_{pd_{2}}}\left(\frac{1}{U+\Delta_{pd_{2}}}\right)^{2}\right$.
 $\quad\left.{}+\frac{1}{2U+\Delta_{pd_{1}}+\Delta_{pd_{2}}}\left(\frac{1}{U+\Delta_{pd_{1}}}+\frac{1}{U+\Delta_{pd_{2}}}\right)^{2}\right),$ (2)
-the $t_{pd_{1,2}}$ represent the hopping between the Ni-$d_{1,2}$ orbitals and the O-$p$ orbital, and the energy gap between the lower Hubbard band of the $d$-orbitals and the O-$p$ orbital is given by: $\Delta_{pd_{1,2}}=\varepsilon_{d_{1,2}}-\varepsilon_{p}$, Note that here the double-counting term is already include in $\varepsilon_{d_{1,2}}$.
-Equation 2 presents the second-order approximation result from perturbation theory, which describes the superexchange between two neighboring Ni-$3d$ orbitals in the atomic limit.
-Our result of $J$ for various $U$ can be found in Table 1.
-One can see that as $U$ increases, the AFM coupling coefficient $J$ decreases.
-For example, at $U=3$ eV, $J_{\perp}=0.325$ eV, and at $U=7$ eV, $J_{\perp}=0.135$ eV.
+where $t_{pd_{1,2}}$ denotes the hopping between the Ni-$d_{1,2}$ and the O-$p$ orbitals.
+The energy difference between the lower Hubbard band of the $d$-orbitals and the O-$p$ orbital is given by: $\Delta_{pd_{1,2}}=\varepsilon_{d_{1,2}}-\varepsilon_{p}$, with the double-counting term included in $\varepsilon_{d_{1,2}}$.
+Equation 2 represents the second-order approximation result from perturbation theory, which describes the superexchange between two neighboring Ni-$3d$ orbitals in the atomic limit.
+Our result of $J$ for various $U$ can be found in Table 1, where one sees that, as expected, the AFM coupling coefficient $J$ decreases monotonically with increasing $U$.
+For instance, the interlayer couplings $J_{\perp}$ drops from 0.325 eV for $U=3$ eV to 0.135 eV for $U=7$ eV.
 Figure 2: The spin-spin correlation function $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ for four neighboring $d$-orbitals is shown in numbers to demonstrate the relative strength of the antiferromagnetic superexchange couplings in the system.
@@
   7   0.135   0.084   62.3%
-In DQMC calculations, the magnetic correlation functions can be exactly measured at high temperature.
-In Figure 2 (a) and (b), we show the magnitudes of the spin correlation function $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ between neighboring $d$- orbitals at $T=0.25$ and half-filling from DQMC, and $\mu=0$ at $T=0.125$ from CDMFT, respectively.
+In DQMC calculations, the magnetic correlation functions can be exactly measured at high temperatures.
+Figure 2 (a) and (b) show the magnitudes of the spin correlation function $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ between neighboring $d$- orbitals at $T=0.25$ and half-filling from DQMC, and $\mu=0$ at $T=0.125$ from CDMFT, respectively.
 Here $S_{ia\alpha}$ is spin operator with site index $i$, layer $a$ and orbital $\alpha$.
 This results depicts the relative strengths of the major exchange couplings in the LNO thin film.
-In Figure 2 (a), at $T=0.25$ and half-filling, comparing to the bulk LNO, the dominant AFM exchange coupling is also the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling, with a value of $\langle S\cdot S\rangle=-0.057(2)$.
+At $T=0.25$ and half-filling, like in bulk LNO, the dominant AFM exchange coupling in films is also the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling , with a value of $\langle S\cdot S\rangle=-0.057(2)$.
 The next strongest AFM exchange comes from IR nearest-neighboring (NN) $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ orbitals, with $\langle S\cdot S\rangle=-0.046(7)$.
-Relative to bulk, this value increases significantly in thin films, reaching 81% of AFM exchange between IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ orbitals.
-The same ratio in bulk LNO it is only 60% [15].
-Moreover, the strengths of IR NN $d_{x^{2}-y^{2}}-d_{3z^{2}-r^{2}}$ and $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ AFM correlations are $\langle S\cdot S\rangle=-0.024(1)$ and $\langle S\cdot S\rangle=-0.017(1)$, respectively.
-These couplings play a minor role due to a weaker hopping between $d_{3z^{2}-r^{2}}$ and $p_{x}/p_{y}$ orbitals in films.
-Notably, the IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ exchange coupling within a unit cell is the weakest of these.
-In Figure 2 (b), the dominant AFM exchange coupling is also the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling from CDMFT approaches when $T=0.125$ and $\mu=0$.
-Notably, the strength of the correlation between IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ orbitals is the weakest of the five magnetic couplings.
-In essence, it is clear that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlation dominates at low temperatures, suggesting that $d_{3z^{2}-r^{2}}$ orbitals are the key components responsible for strong-correlation-driven Superconductivity.
+Comparing to that in bulk, the in-plane coupling becomes significant in thin films, reaching 81% of the AFM exchange between IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ orbitals.
+The same ratio in bulk LNO it is only about 60% [15].
+On the other hand, the strengths of IR NN $d_{x^{2}-y^{2}}-d_{3z^{2}-r^{2}}$ and $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ AFM correlations are $\langle S\cdot S\rangle=-0.024(1)$ and $\langle S\cdot S\rangle=-0.017(1)$, respectively.
+These couplings are relatively small due to a weak hopping between $d_{3z^{2}-r^{2}}$ and $p_{x}/p_{y}$ orbitals.
+Notably, the IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ exchange coupling within a unit cell is the weakest among these magnetic couplings.
+In Figure 2 (b), CDMFT results at $T=0.125$ and $\mu=0$ also suggests that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ coupling is predominant in the system, which has a value of $\langle S\cdot S\rangle=-0.168(4)$.
+Notably, in CDMFT at $\mu=0$, the strength of the correlation between IT $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ orbitals is again the weakest of the five magnetic couplings.
+In essence, here it is clear that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ magnetic correlation dominates at different $\mu$ (or fillings) , suggesting that $d_{3z^{2}-r^{2}}$ orbitals should be the key components responsible for the correlation-driven Superconductivity.
 Figure 3: $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ between pairs of nearest-neighboring (NN) d-orbitals as a function of temperature $T$ at pristine configuration, as obtained from DQMC simulations.
 The squares indicate results for the intra-layer (IR) $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ correlations, while the circles show the results for the inter-layer (IT) $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlations.
-Figure 3 shows the temperature dependence of AFM correlations $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ obtained from DQMC simulations at $\mu=0$.
-Compared with the 11-band model of bulk LNO.
-It is observed that the AFM correlations of bulk LNO are stronger than those in the thin film, for both IR and IT interactions.
-This could provide an explanation for the reduced transition temperature $T_{c}$ in the thin film, caused by the weaker AFM superexchange compared to bulk LNO.
-In film, the reduction in the magnetic correlations between the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ orbitals is much more prominent than the reduction in IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component.
-The latter is almost indistinguishable between film and bulk.
-Furthermore, it can be seen that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ magnetic correlation in thin film is not overly stronger than that of IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component, contrasting with the situation in bulk, where the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlation is much stronger than the IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ correlations [15].
-This results imply that both the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ component and the IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component of AFM correlations may play crucial roles in the observed superconductivity in the LNO thin film at ambient pressure.
+Figure 3 presents the temperature dependence of AFM correlations $\langle S_{ia\alpha}\cdot S_{jb\beta}\rangle$ obtained from DQMC simulations at $\mu=0$.
+It is clear that bulk LNO system exhibits stronger AFM correlations than those in the thin film for both IR and IT components.
+This overall suppression of magnetic couplings provides a plausible explanation for the reduced superconducting transition temperature $T_{c}$ in thin films.
+A key distinction in the suppression of magnetic couplings should be noticed: the reduction of magnetic correlations for the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ component is much more prominent than that for the IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component.
+The latter appears almost identical between film and bulk systems.
+Furthermore, it can be seen that the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ magnetic correlation in thin film is not significantly stronger than the IR NN $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component.
+This contrasts sharply with the bulk case, where the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ correlation is much stronger than its IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ counterpart[15].
+These results suggests that both the IT $d_{3z^{2}-r^{2}}-d_{3z^{2}-r^{2}}$ and the IR $d_{x^{2}-y^{2}}-d_{x^{2}-y^{2}}$ component of AFM correlations may play crucial roles in the ambient-pressure superconductivity observed in LNO thin films.
 Figure 4: Charge transfer properties and charge carrier evolution of the La3Ni2O7 thin film.
@@
 DQMC and CDMFT result of hole concentration $n_{h}$ as a function of hole chemical potential $\mu_{h}$.
-An inflection point suggests that a charge transfer gap opens at half-filling when the hole chemical potential $\mu_{h}$ approaches $\mu_{h}\sim-1.55$.
+An inflection point suggests the opening of a charge transfer gap at half-filling when the hole chemical potential $\mu_{h}$ approaches $\mu_{h}\sim-1.55$.
+Here $T=0.3$.
 (b).
@@
 Here, $\alpha$ denotes distinct $\alpha$-orbitals and charge concentration variation $\Delta n_{e/h}^{\alpha}\equiv n_{e/h}(\mu)-n_{e/h}(\mu_{e}=1.55)$.
-The results for (c), (d) and (e) above are from DQMC simulations at $T=0.3$.
-Charge transfer properties.- We now focus on the charge transfer properties in 11-band Hubbard model, Figure 4 illustrates the vary of the hole/electron concentration changes in different orbitals as functions of chemical potentials $\mu$ and Hund’s coupling $J_{H}$.
-We observe an inflection point at $n_{h}=1$ ($\mu_{h}\sim-1.55$) with an almost flat compressibility ($\partial n_{h}/\partial\mu_{h}$) in Figure 4 (a), which implies there could an opening charge gap at half-filling.
-Note that here due to the escalated temperature we use, the $\mu-n_{h}$ curve is not fully flat, denoting a not vanishing compressibility.
-Notably, we hypothesize that this charge gap is smaller compared to that of bulk LNO due to the significant impact of chemical potential on the charge carrier concentration.
-We also present doping-induced, orbital-resolved distribution of doped charge carriers in Figure 4 (e).
-As the hole chemical potential $\mu_{h}\equiv-\mu$ and electron chemical potential $\mu_{e}\equiv\mu$ are tuned from half-filling ($\mu_{h}=-\mu_{e}\approx-1.55$), holes/electrons to populate distinct Ni-$d$ and O-$p$ orbitals.
-This process, quantified by $\Delta n_{h/e}^{\alpha}$, is shown in Figure 4 (e).
-Consistent with bulk LNO [15, 14, 71], doped holes predominantly occupy O-$p$ orbitals (Right of Figure 4 (e)), while doped electrons primarily occupy Ni-$d$ orbitals (Left of Figure 4 (e)).
-This confirms the charge-transfer nature of the insulating state in this system at the half-filling state[60].
-Moreover, we observe that the $d_{x^{2}-y^{2}}$ orbital contain a higher proportion of doped holes than the $d_{3z^{2}-r^{2}}$ orbital in the case of hole doping ($\mu_{h}>-1.55$), as seen in Figure 4 (e).
-This reflects the stronger correlation of the $d_{3z^{2}-r^{2}}$ orbital compared to the $d_{x^{2}-y^{2}}$ orbital under hole doping, as evidenced by the significant impairment of superexchange coupling in cuprates caused by hole doping [62, 61].
-To shed light on the effective low-energy model study that neglect oxygen degree of freedom, we group the Ni-$3d$ and O-$p$ orbitals into two classes: the out-of-plane (OP) correlated orbitals, including $d_{3z^{2}-r^{2}}$, $p_{z}$ and $p_{z}^{\prime}$ and the in-plane (IP) correlated orbitals, comprising $d_{x^{2}-y^{2}}$ and $p_{x}/p_{y}$ orbitals.
-And analyze their carrier concentration under hole/electron doping.
-The results, shown in Figure 4 (d), indicated that the additional holes doped into the LNO film are distributed nearly equally between the IP and OP orbitals, with a slight excess of holes in the IP orbitals.
-In the case of electron doping, the ratio of doped electrons in the IP to the OP orbitals is approximately 3:1 until the system reaches half-filling.
-This contrasts with the 1:1 ratio observed in hole doping, which suggests an asymmetric doping-dependent superconducting phase diagram.
-In experiments, Sr-doping introduce extra holes into the system [66], while possible Ce-doping can introduce extra electrons into the pristine system.
-In comparison to bulk LNO, we find that under the same chemical potential variation, the doping variable of LNO film is approximately twice that of bulk LNO [15].
-This suggests that LNO films exhibit enhanced charge transfer capabilities compared to the bulk LNO under in-plane biaxial compression, meaning the LNO thin film possesses a smaller charge transfer gap $\Delta$ [72].
-This finding aligns with recent spectroscopic experimental results [73], which propose that two types of Zhang–Rice singlets exist in the LNO film and that the fundamental classification of in-plane charge transfer properties is identical to that of the cuprate superconductors.
-Furthermore, it is discovered that the presence of Hund’s coupling $J_{H}$ in LNO films under half-filling strongly suppresses orbital polarization, thereby stabilizing a high-spin $S=1$ state.
-This phenomenon is illustrated by the gray and red dashed lines in Figure 4 (b), which correspond to the Hund’s first law.
-In contrast, in the pristine state, the $J_{H}$ coupling is significantly localized, and holes still preferentially occupy the in-plane orbitals, as indicated by the gray and red solid lines in Figure 4 (b).
-Discussion and conclusion.- The observation of superconductivity in La3Ni2O7 thin films at ambient pressure [49, 50] has spurred proposals of several effective interacting models to investigate the pairing symmetry [74, 75, 51, 52], building on insights from ab initio calculations [47, 48, 54, 74, 76].
-These effective models primarily focus on parameterizing the magnetic correlations between Ni $d$-orbitals through direct exchange couplings, while the contribution of oxygen degrees of freedom is neglected.
-In contrast our 11-band Hubbard model, which incorporates both Ni-$3d$ and O-$2p$ orbitals, provides a more comprehensive picture by quantifying the relative strengths of superexchange couplings between distinct Ni-$3d$ orbitals.
-Crucially, we identify two spin-singlet bands exhibiting markedly different hole concentrations ($n_{h}$) and correlation strengths.
-We observe the significant reduction of IT magnetic exchange coupling, which makes the IR and IT coupling strengths comparable, in sharp contrast to the IT-dominated bulk material.
-Furthermore, we observe a pronounced asymmetry in charge distribution upon doping.
-It should be noted that some studies suggest Hund’s coupling $J_{H}$ can transfer antiferromagnetic correlation from $d_{3z^{2}-r^{2}}$ orbital to $d_{x^{2}-y^{2}}$ orbital [21, 20, 19, 77].
-Based on our results, we find that a large $J_{H}$ does indeed enhance the inter-layer $d_{x^{2}-y^{2}}$ correlation.
-However, the large $J_{H}$ does not significantly impact the carriers distribution in the $e_{g}$ orbitals.
-Our analysis of both spin-spin correlations and charge-transfer properties consistently indicates that the inter-layer $Ni-3d_{3z^{2}-r^{2}}$ orbital antiferromagnetic exchange interaction remains an important factor that primarily determines the nature of the bilayer nickelate superconducting pairing.
-And the contribution of more itinerant Ni-$d_{x^{2}-y^{2}}$ orbital in to both strange-metal (SM) and superconducting (SC) states remains complex and requires further investigations.
+Results in (c), (d) and (e) above are from DQMC simulations at $T=0.3$.
+Charge transfer properties.- We now examine the charge transfer properties within the 11-band Hubbard model.
+Figure 4 illustrates the evolution of the hole/electron concentrations across different orbitals as functions of chemical potentials $\mu$ and Hund’s coupling $J_{H}$.
+In Figure 4(a), hole concentration $n_{h}$ is shown as a function of $\mu$, where an inflection point near $n_{h}=1$ ($\mu_{h}\sim-1.55$) with strongly reduced compressibility ($\partial n_{h}/\partial\mu_{h}$) can be seen, suggesting the opening of a charge gap at half-filling ($n_{h}=1$).
+For the temperatures we have accessed ($T>0.125$), the $\mu-n_{h}$ curve is not completely flat, indicating that the chager gap is not fully open.
+Notably, the reduced flatness of the $\mu-n_{h}$ curve around $n_{h}=1$ for the film (compared to the bulk at this temperature [15]) should point to a smaller charge transfer gap in the system.
+Figure 4 (e) presents orbital-resolved distribution of doped charge carriers.
+As the hole chemical potential $\mu_{h}\equiv-\mu$ or electron chemical potential $\mu_{e}\equiv\mu$ are tuned from half-filling, doped carriers enter into different orbitals.
+Specifically, we see that doped holes predominantly occupy O-$p$ orbitals ( Blue and Green lines in the right panel of Figure 4 (e)), while doped electrons primarily occupy Ni-$d$ orbitals ( Red and black curves in the left panel of Figure 4 (e)).
+This observation confirms the charge-transfer nature of the correlated states in the system[62, 63].
+Moreover, we observe that the $d_{x^{2}-y^{2}}$ orbital contains a higher concentration of doped holes than the $d_{3z^{2}-r^{2}}$ orbital in the case of hole doping ($\mu_{h}>-1.55$).
+This suggests that the $d_{3z^{2}-r^{2}}$ orbitals are more correlated comparing to the $d_{x^{2}-y^{2}}$ orbital under hole doping.
+To shed light on low-energy effective models integrating out oxygen degrees of freedom, we group the Ni-$3d$ and O-$2p$ orbitals into two categories: (i) the out-of-plane (OP) correlated orbitals, including $d_{3z^{2}-r^{2}}$, $p_{z}$ and $p_{z}^{\prime}$ and (ii) in-plane (IP) correlated orbitals comprising $d_{x^{2}-y^{2}}$ and $p_{x}/p_{y}$ orbitals.
+As shown in Figure 4 (d), additional holes doped into LNO film distribute nearly equally between IP and OP orbitals.
+In contrast, for electron doping around $\mu_{e}\sim 0$, the doped carriers favor IP orbitals over OP orbitals by a ratio of approximately 3:1.
+This stark asymmetry between electron and hole doping (1:1 vs.
+3:1) suggests a dopant-type dependent superconducting phase diagram.
+This is particularly relevant to experiment, where Sr-doping introduces holes [68], while Thorium (Th) doping could introduce electrons.
+The data in Figure 4(d) also show a significant difference in carrier response: under identical chemical potential changes, the hole variation in LNO films is about twofold greater than in the bulk [15].
+This indicates that LNO thin films possess enhanced charge transfer capabilities and, consequently, a smaller charge transfer gap $\Delta$ [74].
+Our conclusion aligns with recent spectroscopic experiments [75], which propose the existence of two types of Zhang–Rice singlets in LNO films and suggest the in-plane charge transfer properties are analogous to those of cuprate superconductors.
+Finally, we find that the presence of Hund’s coupling $J_{H}$ at half-filling strongly in general suppresses orbital polarization, hence tends to stabilize a high-spin $S=1$ state, as illustrated by the dashed lines in Figure 4 (b), which follows the Hund’s first law.
+However, for the pristine configuration ($\mu=0$), the influence of the $J_{H}$ coupling is significantly weaken.
+In this case, holes retain a preferential occupation of the in-plane orbitals, as indicated by the solid lines in Figure 4 (b).
+Discussion and conclusion.- The observation of superconductivity in La3Ni2O7 thin films at ambient pressure [51, 52] has spurred proposals of several effective interacting models for investigating various physical properties [76, 77, 53, 54], building on insights from ab initio calculations [49, 50, 56, 76, 78].
+These effective models primarily focus on parameterizing the magnetic correlations based on direct exchange couplings between nickel’s $d$-orbitals, while the contribution of oxygen degrees of freedom is neglected.
+In this work, the 11-band dpd-p Hubbard model, which incorporates both Ni-$3d$ and O-$2p$ orbitals, provides a more comprehensive picture for understanding the relative strengths of superexchange couplings and distributions of charge carriers in La3Ni2O7 thin films.
+Crucially, we resolve properties of in-plane and out-of-plane bands exhibiting remarkable different hole concentrations ($n_{h}$) and correlation strengths.
+We find that IR and IT magnetic coupling strengths are more or less comparable, in sharp contrast to the IT coupling - dominated bulk materials.
+Furthermore, we observe a pronounced asymmetry in charge distribution upon hole/electron doping.
+In summary, our unified analysis on both magnetic correlations and charge-transfer properties consistently identifies the interlayer $Ni-3d_{3z^{2}-r^{2}}$ orbital couplings as the dominant factor determining the nature of pairing in the bilayer nickelate films.
+The contribution of the more itinerant Ni-$d_{x^{2}-y^{2}}$ orbitals in superconducting or normal state, however, appear more subtle and complex, warranting further dedicated investigations.
 Acknowledgements.- We thank the useful discussion with Zhihui Luo and Wenyuan Qiu.
-This project was supported by NSFC-12494591, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).
+This project was supported by NSFC-12494591, NSFC-12494594, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).